OpenStudy (anonymous):

solve by using the quadratic formula, x(x-3) = 4

OpenStudy (anonymous):

x^2-3x-4=0 D=9+16=25 x1=(3+5)/2=4 x2=(3-5)/2=-1

OpenStudy (anonymous):

man you are on a roll, haha, thanks again